OpenStudy (anonymous):

Ahmed is taking orders for lunch. A slice of pizza cost $2.50 and a chicken sandwich cost $3.50. He collects $21.50 for the group of 7. let x= number of pizza orders. Let y= number of chicken orders. Assuming each person only orders one item. how many people order the chicken sandwich?

OpenStudy (mathstudent55):

What is x + y = ?

OpenStudy (anonymous):

3.50x+2.50y=21.50?

OpenStudy (mathstudent55):

You've got the right idea, but you switched the prices. x = number of pizza slices. Slice costs $2.50 y = number of chicken sandwiches. Sandwich costs $3.50 x goes with 2.50 and y goes with 3.50

OpenStudy (anonymous):

2.50x+3.50y=21.50, so I solve for y?

OpenStudy (mathstudent55):

This is one equation. You have two variables, so you need a second equation. That's why I asked what is x plus y equal to?

OpenStudy (anonymous):

Oh. So. 7 x $2.50 = $21.50 1 x $3.50 = $3.50

OpenStudy (anonymous):

I'm confused.

OpenStudy (mathstudent55):

You wrote the equation below, and it is good. 2.50x+3.50y=21.50 Now you need one more equation that connects x and y. There is a piece of information the problem gave that will allow you to write a second equation that also has x and y.

OpenStudy (mathstudent55):

There are 7 people placing orders. Each person only orders one order, either a slice or a sandwich. So we know that only 7 orders will be placed. There will be x orders of pizza and y orders of chicken sandwich. The total number of orders is 7, so x + y = 7

OpenStudy (mathstudent55):

Now you have two equations with two unkowns. Now you can solve the system of equations for x and y.

OpenStudy (anonymous):

y=4 x=3

OpenStudy (anonymous):

right?

OpenStudy (mathstudent55):

2.50x+3.50y=21.50 x + y = 7 This system of equation needs to be solved. Take the second equation and solve it for x: x = 7 - y Now insert what x is equal to, 7 - y, into x of the first equation: 2.5(7 - y) + 3.5y = 21.5 17.5 - 2.5y + 3.5y = 21.5 y + 17.5 = 21.5 y = 4 x + y = 7 x + 4 = 7 x = 3 Answer: 3 slices & 4 sandwiches

OpenStudy (mathstudent55):

You are correct